# Of Mice and Machines: A Comparison of Learning Between Real World Mice and RL Agents

### v1

@@ 1 Introduction
Mathematical foundations for reinforcement learning (RL) emerged through several landmark contributions: Samuel’s early work on learning systems, Bellman’s dynamic programming principles, and later Sutton’s temporal difference learning (Samuel, 1959; Bellman, 1966; Sutton, 1988). While these seminal advances have enabled powerful decision-making agents, they arose primarily from computational principles rather than biological experimentation.  [^1] While Sutton et al. (1998) has some biological inspiration, there is a clear distinction to be drawn between biologically inspired and derived from measured biological data. We investigate the latter viewpoint in this volume. Although subsequent research has attempted to draw parallels between RL algorithms and biological learning mechanisms (Pozzi et al., 2018; Neftci & Averbeck, 2019; Tan et al., 2023), suggesting the biological plausibility of methods like Q-learning, a fundamental question remains: To what extent do RL agents truly capture the decision-making characteristics of biological organisms?
@@
To investigate this question, we designed a controlled experimental environment where mice navigate a complex space while avoiding a robotic predator-like autonomous agent (hereafter referred to as “predator” for brevity). A simulation environment replicates this with the prey controlled by an RL algorithm. The predator is a reactive agent in both the experiment and simulation. Through careful behavioral analysis, offline RL modeling (Levine et al., 2020) of mouse behavior, and Exploratory Data Analysis over visitation density graphs, we identified several striking disparities between biological and artificial agents. Most notably, RL agents demonstrate a remarkable lack of self-preservation instinct, often choosing marginally more efficient paths that bring them dangerously close to the predator. In contrast, biological agents exhibit sophisticated risk-assessment behaviors, with mice spending over 50% of their time gathering environmental information and evaluating predator positions before movement.
@@
Inspired by these observations, we developed two novel mechanisms to bridge the behavioral gap. First, we introduce a modified replay buffer mechanism that amplifies and frequently resamples near-death experiences during training, mimicking post-traumatic stress responses observed in biological systems. This enhancement enables agents to develop more appropriate risk-avoidance behaviors. Second, we propose a fundamental modification to the TD-learning framework incorporating action uncertainty through Q-value variance estimation. This addition creates naturally risk-averse agents that better reflect the cautious decision-making patterns observed in biological systems. Empirically, These mechanisms enhance the behavioral similarity between artificial and biological agents, increasing the visitation pattern overlap with mice from 20.9% to 86.1%. The enhanced agents exhibit conservative movement patterns with approximately 45% less distance traveled during initial environment entry, closely matching the cautious behavior of real mice.
@@
  • Results show that incorporating this mechanism leads to trajectories that are closer to biological behavior. This is true even when compared to other surprise-avoiding objectives from the literature such as SMiRL. (Berseth et al., 2019)
@@ 2 Experiment Setup
This study assessed strategic evasion behaviors in mice through a task approximating predator-prey interactions within a hexagonal arena, namely Cellworld (Lai et al., 2024). The physical setup included a custom robotic predator-like threat that pursued the mice, simulating predatory threat through aversive air puffs, while the mice attempted to reach a designated goal area without being “captured” (within 27.5 cm or 0.1 units of the robot). The arena was mapped with a hexagonal grid of 331 magnetic cells, each 11 cm apart, where obstacles were strategically placed to encourage adaptive evasion strategies.
@@
For a direct comparison with the reinforcement learning (RL) model, we created a simulated environment, Cellworld Gymnasium, building upon the Gymnasium API (Towers et al., 2024) and Pygame, which closely mirrors the physical experiment setup. This simulated environment replicates the hexagonal grid layout of the real arena and incorporates the same reward structure: the virtual “mouse” receives a reward of +1 upon reaching the goal without capture and a penalty of -1 if it is “puffed” by the simulated predator-like robotic threat. [^2] The equal magnitude (+1/-1) for rewards and penalties was chosen to reflect that both water reward and air puff are relatively mild stimuli. To further ensure fidelity, both the mouse and the predator in the RL environment move at speeds proportionally scaled to the average speeds observed in the physical trials. Additionally, the RL agent has partial observations, detecting the predator’s location only when the predator is within its field of view. All distances within the arena are normalized such that 1 unit corresponds to the diameter of the hexagonal arena.
@@ 3.1.1 Model-Free Methods
Deep Q-Network (DQN): Utilizes a neural network to approximate the action-value function $Q(s,a)$. The update rule is based on the Bellman equation (Hester et al., 2018):
@@ 4 Comparison Between Mouse Behavior and RL Agent Behavior
By directly comparing mouse behavior with RL agents in the Cellworld setting, we aim to characterize key behavioral differences in their threat evasion strategies. Specifically, we analyze trajectory patterns to quantify wall-following tendencies, examine state density distributions to understand waiting behaviors near obstacles and starting points and compare action sequences before and after predator detection. Through these analyses, we seek to uncover the fundamental differences between mice and RL agents in risk management in predator-prey scenarios.
We begin by training TDMPC-2 on the Cellworld Gymnasium environment and comparing the behaviors generated to those of the mice. On a superficial level, the performances of the mice and TDMPC-2 are similar, with the mice achieving an 86% success rate across environments and the RL algorithms achieving about 80%. However, when we examine the exact nature of these successes, we see there is a dramatic divergence between the two agents.
@@
As shown in Figure 3, mice trajectories exhibit distinct patterns concentrated along walls (thigmotaxis, commonly observed in many animals and corresponding to measures of anxiety (Mugan & MacIver, 2020)) and obstacles, with approximately 70% of trajectories following wall-following paths. [^3] Defined as trajectories where over 70% of points are within 0.1 units (10% of the diameter of the hexagonal arena) of boundaries. In contrast, RL agents’ trajectories show direct, goal-oriented paths with minimal environmental interaction, with 0% of trajectories hugging walls. The state-visitation density of RL agents is heavily concentrated along these direct routes, showing little of the preemptive caution or environmental awareness displayed by the mice.
@@
Figure 6 highlights key behavioral differences between TDMPC-2 and mice. In distance changes (top row), TDMPC-2 shows excessive movement both before (-0.2 to -0.4) and after (-0.2 to -0.35) predator detection, while mice maintain conservative patterns (-0.05 to -0.15), reflecting natural cautiousness. Spatially (bottom row), mice trajectories exhibit structured, radial exploration patterns, whereas TDMPC-2’s movements appear scattered and disorganized, lacking the strategic positioning characteristic of natural behavior.
Another key difference lies in learning efficiency after negative encounters. RL agents typically require multiple “puffs” before adapting their behavior to avoid the predator, while mice exhibit immediate and persistent avoidance after a single encounter. Of course, this disparity is natural—mice have millions of years of evolutionary risk avoidance encoded in their genetics, while neural networks begin learning from scratch. However, this gap highlights an important challenge in artificial intelligence: how to develop RL systems that can more readily incorporate risk-avoidance behaviors without requiring extensive negative experiences.
@@ 5.1 Trauma-Inspired Safety Memory in RL Agents
Can we teach RL agents to value their “lives” and fear predators?
A common approach is Prioritized Experience Replay (PER) (Schaul et al., 2015). Our experiments show that PER brings some benefits, but doesn’t fully solve the challenges in our scenario. After 50,000 training steps, PER yields an average reward improvement of about 5, yet the agents’ behaviors change only marginally. In other words, PER speeds up learning but has little effect on how the agents actually behave. This shortcoming inspired our new extensions.
@@
Figure 7: A performance comparison of Selective Memory settings over the first 5000 steps reveals that a 50% setting yields the highest efficiency. The experiments were conducted using 5 different random seeds.
Selective Memory Sampling: The buffer prioritizes traumatic experiences (events with negative rewards), ensuring they make up 50% of each training batch. This mimics trauma-like memory recall, helping the agent focus on learning from dangerous situations. Experiments (Figure 7) indicate that 50% (ranging from 0% to 100% in increments of 25%) is the optimal percentage for this buffer. [^4] This rebalancing mimics rapid early learning in mice while primarily benefiting early training stages; effects diminish or can become counterproductive once the agent accumulates diverse experiences.
Memory Amplification: Negative experiences are amplified by a factor of 200 during sampling [^5] We conducted experiments with different amplification factors: smaller values showed limited behavioral changes, while factors above 200 slowed down training without providing substantial benefits., reflecting the heightened emotional intensity of trauma in biological systems, which leads to stronger behavioral changes.
Figure 8: Density plots of 50% Selective Memory on TDMPC-2. Left: The baseline matches the standard TDMPC-2’s core behavior (Figure 4 left) with minor tail improvements; Right: Amplification (200) increases path diversity and wall-following behavior.
Our computational experimental results demonstrate that the TISB induces more naturalistic behaviors in RL agents. Figure 8 illustrates broader exploration and wall-following patterns similar to those observed in mice. However, the agents still lack the characteristic waiting behaviors seen in mice, as evidenced by the significantly lower density at the entrance compared to mice. This observation motivates our next approach.
@@ 5.2 Learning to Wait Through Variance-Penalized Temporal Difference (TD) Learning
Figure 9: Density plots: TDMPC-2 (left), SMiRL without TISB (middle), and SMiRL with TISB (right). SMiRL may negate TISB’s effect, as the main density distribution shows little change.
Unfortunately, incorporating SMiRL into TDMPC-2 did not have the desired effect of making RL agents more conservative and cautious. As shown in Figure 9, while SMiRL shows some increased density in the central area compared to the baseline TDMPC-2, suggesting slightly more diverse path choices, the core trajectory pattern remains largely unchanged, as evidenced by the similar high-density regions (bright pink areas) across all conditions. Most importantly, SMiRL failed to demonstrate the crucial waiting behavior at the entrance, which was our primary goal. Furthermore, when combined with TISB, SMiRL appears to counteract TISB’s path-finding benefits that were previously observed in TDMPC-2, resulting in behavior that remains closer to the baseline approach rather than achieving the fundamental strategy shift we were aiming for.
@@
Figure 10: Final states visitation pattern comparison of original TDMPC-2 (left), to our VP-TDMPC2 (middle: trained with Memory Amplification and Variance-Penalized TD learning) and Mouse (right).
Figure 11: Top row: Mean changes in goal-agent distance before and after predator detection (left, middle), and initial mean movement distances (right). Shaded areas show standard deviation. Bottom row: Spatial distribution of first three steps for each agent, with black stars indicating mean positions. (1 unit = the diameter of the hexagonal arena)
@@
[…]
@@
Notably, VP-TDMPC-2 exhibits geometric patterns and balanced coverage resembling mice’s systematic exploration, though it lacks the extreme thigmotaxis (wall-following with minimal distance throughout the entire trail) observed in real mice. For a closer examination of the differences between mice behavior and RL, we analyzed the initial part of the trajectory from the start gate.
@@
Episode length also differ substantially (Figure 13 in Appendix). Standard RL agents completed the task in an average of 7.09 steps, while VP-TDMPC-2 averaged 13.89 steps, closely aligning with the 14.04-step average observed in mice. Longer episodes may reflect more cautious or deliberate planning under risk.
@@ 5.3 Challenges in Closing the Mice and Machines Gap
Our hypothesis is that this behavioral difference arises from the structure of the RL state space. After thoroughly exploring the outer wall during training, the agent no longer needs to maintain constant wall contact for safety. Upon reaching the top of the arena, it can confidently conclude that the task is solved. Additionally, the original TDMPC2 agent relied almost entirely on the central regions. Although our improved RL agent has learned to exhibit some cautious, mouse-like behavior, it has also learned that the central area offers more efficient paths to the goal when safety is ensured. In contrast, mice continue to follow the wall regardless of task mastery, likely due to rodents’ innate behavioral priors (Champagne et al., 2010).
@@ 7 Related Work
The discovery of mirror neurons, which exhibit similar activation patterns during both action execution and observation (Buccino et al., 2004), has provided a biological foundation for imitation learning, influencing both robotic systems (Zahra et al., 2022) and third-person imitation learning algorithms in RL (Stadie et al., 2017). Recent advances in distributional reinforcement learning suggest that the brain encodes not merely the mean expected reward, but rather the complete distribution of possible rewards (Muller et al., 2024).

### v3

@@ 1 Introduction
Mathematical foundations for reinforcement learning (RL) emerged through several landmark contributions: Samuel’s early work on learning systems, Bellman’s dynamic programming principles, and later Sutton’s temporal difference learning (Samuel, 1959; Bellman, 1966; Sutton, 1988). While these seminal advances have enabled powerful decision-making agents, they arose primarily from computational principles [^1] While Sutton et al. (1998) has some biological inspiration, there is a clear distinction to be drawn between biologically inspired and derived from measured biological data. We investigate the latter viewpoint in this volume. rather than biological experimentation. Although subsequent research has attempted to draw parallels between RL algorithms and biological learning mechanisms (Pozzi et al., 2018; Neftci & Averbeck, 2019; Tan et al., 2023), suggesting the biological plausibility of methods like Q-learning, a fundamental question remains: To what extent do RL agents truly capture the decision-making characteristics of biological organisms?
@@
To investigate this question, we designed a controlled experimental environment where mice navigate a complex space while avoiding a robotic predator-like autonomous agent (hereafter referred to as “predator” for brevity). A simulation environment replicates this with the prey controlled by an RL algorithm. The predator is a reactive agent in both the experiment and simulation. Through careful behavioral analysis, offline RL modeling (Levine et al., 2020) of mouse behavior, and Exploratory Data Analysis over visitation density graphs, we identified several striking disparities between biological and artificial agents. Most notably, RL agents demonstrate a remarkable lack of self-preservation instinct, often choosing marginally more efficient paths that bring them dangerously close to the predator. In contrast, biological agents exhibit sophisticated risk-assessment behaviors, with mice spending a significant amount of time gathering environmental information and evaluating predator positions before movement.
@@
Inspired by these observations, we developed two novel mechanisms to bridge the behavioral gap. First, we introduce a modified replay buffer mechanism that amplifies and frequently resamples near-death experiences during training, mimicking post-traumatic stress responses observed in biological systems. This enhancement enables agents to develop more appropriate risk-avoidance behaviors. Second, we propose a fundamental modification to the TD-learning framework incorporating action uncertainty through Q-value variance estimation. These additions create naturally risk-averse agents that better reflect the cautious decision-making patterns observed in biological systems. Empirically, these mechanisms enhance the behavioral similarity between artificial and biological agents, increasing the visitation pattern overlap with mice from 20.9% to 86.1%. The enhanced agents exhibit conservative movement patterns with approximately 45% less distance traveled during initial environment entry, closely matching the cautious behavior of real mice.
@@
  • Results show that incorporating these mechanisms lead to trajectories that are closer to biological behavior. This is true even when compared to other surprise-avoiding objectives from the literature such as SMiRL. (Berseth et al., 2019)
@@ 2 Experiment Setup
This study assessed strategic evasion behaviors in mice through a task approximating predator-prey interactions within a hexagonal arena, namely “cellworld” (Lai et al., 2024). The physical setup included a custom robotic predator-like threat that pursued the mice, simulating predatory threat through aversive air blasts, while the mice attempted to reach a designated reward area without being “captured” (within 27.5 cm or 0.1 units of the robot). The arena was mapped with a hexagonal grid of 331 magnetic cells, each 11 cm apart, where obstacles were strategically placed to encourage adaptive evasion strategies(Mugan & MacIver, 2020).
@@
For a direct comparison with the reinforcement learning (RL) model, we created a simulated environment, “cellworld gymnasium”, building upon the Gymnasium API (Towers et al., 2024) and Pygame, which closely mirrors the physical experiment setup. This simulated environment replicates the hexagonal grid layout of the real arena and incorporates the same reward structure: the virtual “mouse” receives a reward of +1 upon reaching the goal without capture and a penalty of -1 if it is “puffed” by the simulated predator-like robotic threat. [^2] The equal magnitude (+1/-1) for rewards and penalties was chosen to reflect that both water reward and air blasts are relatively mild stimuli. To further ensure fidelity, both the mouse and the predator in the RL environment move at speeds proportionally scaled to the average speeds observed in the physical trials. Additionally, the RL agent has partial observations, detecting the predator’s location only when the predator is within its field of view. All distances within the arena are normalized such that 1 unit corresponds to the diameter of the hexagonal arena.
@@ 3.1.1 Model-Free Methods
Deep Q-Network (DQN): It utilizes a neural network to approximate the action-value function $Q(s,a)$. The update rule is based on the Bellman equation (Hester et al., 2018):
@@ 4 Comparison Between Mouse Behavior and RL Agent Behavior
By directly comparing mouse behavior with RL agents in the “cellworld” setting, we aim to characterize key behavioral differences in their threat evasion strategies. Specifically, we analyze trajectory patterns to quantify wall-following tendencies, examine state density distributions to understand waiting behaviors near obstacles and starting points and compare action sequences before and after predator detection. Through these analyses, we seek to uncover the fundamental differences between mice and RL agents in risk management during predator-prey interactions.
We begin by training TDMPC-2 on the “cellworld gymnasium” environment and comparing the behaviors generated to those of the mice. On a superficial level, the performances of the mice and TDMPC-2 are similar, with the mice achieving an 86% success rate [^3] Proportion of runs in which the agent reached the goal without being captured. across environments and the RL algorithms achieving about 80%. However, when we examine the exact nature of these successes, we see there is a dramatic divergence between the two agents.
@@
As shown in Figure 3, mice trajectories exhibit distinct patterns concentrated along walls (thigmotaxis, commonly observed in many animals and corresponding to measures of anxiety (Mugan & MacIver, 2020) and obstacles, with approximately 70% of trajectories following wall-following paths. [^4] Defined as trajectories where over 70% of states are within 0.1 units (10% of the diameter of the hexagonal arena) of the wall. In contrast, RL agents’ trajectories show direct, goal-oriented paths with minimal environmental interaction, with 0% of trajectories hugging walls. The state-visitation density of RL agents is heavily concentrated along these direct routes, showing little of the preemptive caution or environmental awareness displayed by the mice.
@@
Figure 6 highlights key behavioral differences between TDMPC-2 and mice. In reduction of distance to goal (top row), TDMPC-2 shows excessive movement both before (-0.2 to -0.4) and after (-0.2 to -0.35) predator detection, while mice maintain conservative patterns (-0.05 to -0.15), reflecting natural cautiousness. Spatially (bottom row), mice trajectories exhibit structured, radial exploration patterns, whereas TDMPC-2’s movements appear scattered and disorganized.
Another key difference lies in learning efficiency after negative encounters. RL agents typically require multiple air blasts before adapting their behavior to avoid the predator, while mice exhibit immediate and persistent avoidance after a single encounter. Of course, this disparity is natural—mice have millions of years of evolutionary risk avoidance encoded in their genetics, while neural networks begin learning from scratch. However, this gap highlights an important challenge in artificial intelligence: how to develop RL systems that can more readily incorporate risk-avoidance behaviors without requiring extensive negative experiences.
@@ 5.1 Trauma-Inspired Safety Memory in RL Agents
Can we teach RL agents to value their own ’lives’ more and have a greater fear of the predator?
A common approach is Prioritized Experience Replay (PER) (Schaul et al., 2015). Our experiments demonstrate that PER provides some benefits, but does not fully address the challenges in our scenario. While PER accelerates learning, it has minimal impact on the agents’ behavioral policies. This limitation motivated the development of our novel extensions.
@@
Figure 7: Experiments over 5000 steps show that 50% negative experiences yield optimal performance (5 random seeds).
Threat Experience Rebalancing: Inspired by how rodents show heightened sensitivity to dangerous situations, we investigate whether a higher proportion of negative experiences is necessary for effective learning. We employ various percentage settings to ensure that each training batch contains a specified proportion of negative transitions during sampling. Our experiments (Figure 7) shows that 50% negative experiences in the training batch yields optimal performance. This finding provides valuable insights into the importance of negative experiences during early learning phases, similar to how young animals develop risk assessment behaviors.
Threat Experience Amplification: The rewards of negative experiences are amplified by a factor of 200 during sampling [^5] We conducted experiments with different amplification factors: smaller values showed limited behavioral changes, while factors above 200 slowed down training without providing substantial benefits., reflecting the heightened emotional intensity of trauma in biological systems, which leads to stronger behavioral changes.
Figure 8: Standard TDMPC-2 with limited exploration patterns; Right: With a Threat Experience Amplification factor of 200, showing increased path diversity and enhanced wall-following behavior.
Our computational experimental results demonstrate that the TISB induces more naturalistic behaviors in RL agents. In our long-term experiments, TISB was implemented primarily through the Threat Experience Amplification mechanism. Figure 8 illustrates broader exploration and wall-following patterns similar to those observed in mice. However, the agents still lack the characteristic waiting behaviors seen in mice, as evidenced by the significantly lower density at the entrance compared to mice. This observation motivates our next approach.
@@ 5.2 Learning to Wait Through Variance-Penalized Temporal Difference (TD) Learning
Figure 9: Density plots: TDMPC-2 (left), SMiRL without TISB (middle), and SMiRL with TISB (right). SMiRL may negate TISB’s effect, as indicated by only minor changes in the main density regions.
Unfortunately, incorporating SMiRL into TDMPC-2 did not have the desired effect of making RL agents more conservative and cautious. As shown in Figure 9, while SMiRL shows some increased density in the central area compared to the baseline TDMPC-2, suggesting slightly more diverse path choices, the core trajectory pattern remains largely unchanged, as evidenced by the similar high-density regions (bright pink areas) across all conditions. Most importantly, SMiRL failed to demonstrate the crucial waiting behavior at the entrance, which was our primary goal. Furthermore, when combined with TISB, SMiRL appears to counteract TISB’s path-finding benefits that were previously observed in TDMPC-2, resulting in only minor changes to the density distribution compared to the more substantial strategy shifts we were aiming for.
@@
Figure 10: Top row: Mean changes in goal-agent distance before and after predator detection (left, middle), and initial mean movement distances (right). Shaded areas show standard deviation. Bottom row: Spatial distribution of first three steps for each agent, with black stars indicating mean positions. The black stars show that mouse data and VP-TDMPC2 maintain more consistent, cautious initial positions, while baseline TDMPC2’s mean positions spread out more rapidly, indicating less conservative early behavior. (1 unit = the diameter of the hexagonal arena)
@@
$\text{SMiRL:}\quad\max_{\pi}\mathbb{E}\left[\sum_{t}\gamma^{t}\left(r_{t}-\alpha_{1}\log p(s_{t}\mid s_{<t})\right)\right]$
$\text{Our approach:}\quad\max_{\pi}\mathbb{E}\left[\sum_{t}\gamma^{t}\left(r_{t}-\alpha_{2}\text{Var}_{a}(Q(s_{t},a))\right)\right]$
@@
Figure 11: Final states visitation pattern comparison of original TDMPC-2 (left), to our VP-TDMPC2 (middle: trained with TISB and Variance-Penalized TD learning) and Mouse (right).
Notably, VP-TDMPC-2 exhibits geometric patterns and balanced coverage resembling mice’s systematic exploration, though it lacks the extreme thigmotaxis (wall-following with minimal distance throughout the entire trial) observed in real mice. For a closer examination of the differences between mice behavior and RL, we analyzed the initial part of the trajectory from the start gate.
@@
Episode length also differed substantially (Figure 13 in Appendix). Standard RL agents completed the task in an average of 7.09 steps, while VP-TDMPC-2 averaged 13.89 steps, closely aligning with the 14.04-step average observed in mice. Longer episodes may reflect more cautious or deliberate planning under risk.
@@ 5.3 Challenges in Closing the Mice and Machines Gap
Our hypothesis is that this behavioral difference arises from the structure of the RL state space. After thoroughly exploring the outer wall during training, the agent no longer needs to maintain constant wall contact for safety. Upon reaching the top area of the arena, it can confidently conclude that the task is solved. Additionally, the original TDMPC2 agent relied almost entirely on the central regions. Although our improved RL agent has learned to exhibit some cautious, mouse-like behavior, it has also learned that the central area offers more efficient paths to the goal when safety is ensured. In contrast, mice continue to follow the wall regardless of task mastery, likely due to rodents’ innate behavioral priors (Champagne et al., 2010).
@@ 7 Related Work
The discovery of mirror neurons, which exhibit similar activation patterns during both action execution and observation (Buccino et al., 2004), has provided a biological foundation for imitation learning, influencing both robotic systems (Zahra et al., 2022) and third-person imitation learning algorithms in RL (Stadie et al., 2017). Building on biological inspiration, researchers have also integrated curiosity-driven learning mechanisms (Stadie et al., 2015; Oudeyer & Smith, 2016) into RL algorithms to enhance exploration and learning efficiency. Recent advances in distributional reinforcement learning suggest that the brain encodes not merely the mean expected reward, but rather the complete distribution of possible rewards (Muller et al., 2024).